OpenStudy (anonymous):

How long would the car take to go from 22 m/s to 30m/s with the same acceleration?

OpenStudy (jennychan12):

I'm gonna use physics in this. sorry. looking for t (time) so use v = v0+at 30 = 22+(9.8)t 9.8t = 7 t = 0.71 s sorry my calculus with word problems is bleh you could try asking someone else for a calculus explanation?

OpenStudy (jennychan12):

even if you aren't in calculus, sorry.

OpenStudy (anonymous):

i was looking for a physics explanation

OpenStudy (jennychan12):

ohh it was in math group sorry.